How much thermal energy is required to heat all of the water in a swimming pool by 1°C if the dimensions are 4 ft deep



by 20 ft wide by 75 ft long? Report your result in megajoules.

1
Expert's answer
2021-11-11T15:43:01-0500

"1ft=0.3048m"

volume of water in the pool "=(75\times 0.3048)(20\times 0.3048)(4\times 0.3048)=169.90m^3"

Density of water "=1000kg\/m^3"

Mass of water "=169.90\times 1000"

Heat energy "=mc \Delta T = (169.90\times 1000)\times 4200\times 1=713.58 \times 10^6J=713.58MJ"